OpenStudy (anonymous):

A train starting from rest attains of 72 km/h in 5mins assuming that acc. is uniform a. The acc. b. the distance traveled by the train for attaining this velocity.

OpenStudy (anonymous):

@Yahoo! CAN U

OpenStudy (anonymous):

yup plzz wait

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

v= u + at

OpenStudy (anonymous):

use this

OpenStudy (anonymous):

u =0 becoz starting from rest

OpenStudy (anonymous):

72 km/h = 20m/s

OpenStudy (anonymous):

5 min = 300s

OpenStudy (anonymous):

now find acceleration

OpenStudy (anonymous):

use the laws of motion (physics)

OpenStudy (shane_b):

Didn't we do this not too long ago?\[a=\frac{Delta V}{\Delta T}=\frac{20m/s}{(5*60s)}=?\]

OpenStudy (anonymous):

20 = 0 + 300a a = 20/300 = 0.066

OpenStudy (anonymous):

m/s^2

OpenStudy (anonymous):

s = ut + 1/2 a t^2 s = 0 + 1/2 * 0.06 * 300 *300 = 2700 m

OpenStudy (anonymous):

yup